I could really use some help with this problem. Thanks! The area of a rectangle is 45 square cm. If the length is 4 cm greater than the width, what are the dimension of the rectangle?
Answer by ankor@dixie-net.com(22740)

The area of a rectangle is 45 square cm. If the length is 4 cm greater than the width, what are the dimension of the rectangle?
:
Write an equation which translates the following sentence:
"the length is 4 cm greater than the width," (the word "is" usually means "=")
|
L = W + 4
:
We know that:
L * W = 45 sq cm
;
Substitute (W + 4) for L, and you have:
(W+4) * W = 45
:
W^2 + 4W = 45
:
W^2 + 4W - 45 = 0; our old friend the quadratic equation
:
Factors easily to:
(W + 9) (W - 5) = 0
:
W = -9
and
W = +5, the positive solution is what we want in the real world
:
:
L = 5 + 4 = 9: Checks out: 9 * 5 = 45